OpenStudy (jasmineflvs):

-p-4-(2p-5)=4+2(p+3) I just need the basic steps please

OpenStudy (anonymous):

do you look for p?

OpenStudy (anonymous):

-p-4-(2p-5)=4+2p+6

OpenStudy (jasmineflvs):

one more step plz sorry not really "that" great in algebra 1

OpenStudy (anonymous):

-p-4-2p-5=4+2p+6

OpenStudy (jasmineflvs):

thx